Lionblaze18:

Help ASAP 2x+y and 4x+10 must equal 90 degrees, what is the messure for x?

snowflake0531:

\[2x+y+(4x+10)=90\\6x+y+10=90\\6x+y=80\\6x=80-y\\x=\frac{80-y}{6}\]

Lionblaze18:

Thank you